# Bank–Laine functions with preassigned number of zeros

### v2

@@ Abstract.
A Bank–Laine function $E(z)$ is written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation $f^{\prime\prime}+A(z)f=0$ (†), where $A(z)$ is an entire function. We show the existence of entire functions $A(z)$ for which the Bank–Laine functions $E=f_{1}f_{2}$ have preassigned number of zeros of three types:
@@
    for a dense set of orders $\rho\in(1,\infty)$, there exists an entire function $A(z)$ of order $\rho$ such that $E=f_{1}f_{2}$ satisfies $\lambda(E)=\rho(E)=\rho$ and $c_{2}r^{\rho}\leq n(r,0,f_{i})\leq c_{1}r^{\rho}$, $i=1,2$, for two positive constants $c_{1}$ and $c_{2}$ and all large $r$;
@@
    let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$, then there exists an entire function $A(z)$ of order $n$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$;
@@
    let $\rho\in(1/2,\infty)$ and $\lambda\in[0,\infty)$, then there exists an entire function $A(z)$ of order $\rho$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda$ and $\lambda(f_{2})=\infty$ and $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other $f_{3}$ linearly independent from $f_{1}$.
The first result completes the construction of Bergweiler and Eremenko while the proof for the second and third results requires new developments of the method of quasiconformal surgery by Bergweiler and Eremenko.
@@ 1. Introduction
(1.1) $f^{\prime\prime}+A(z)f=0,$
where $A(z)$ is an entire function, under the framework of Nevanlinna theory. See [9] for the standard notation and fundamental results of Nevanlinna theory. Recall that, for a nonconstant entire function $f$, the order $\rho(f)$ of $f$ is defined as
@@
where $m(z)$ is the multiplicity of the zero $z$ and $n(r,0,f)=\sum_{|z|\leq r}m(z)$ denotes the number of zeros of $f$ in the disc $\{z:|z|\leq r\}$. It is elementary that all nontrivial solutions of equation (1.1) are entire functions. When $A(z)$ is a transcendental entire function of finite order, an application of the lemma on the logarithmic derivative [9, Section 2.2] easily yields that all nontrivial solutions of (1.1) satisfy $\rho(f)=\infty$. However, there are examples of $A(z)$ (say $A(z)=e^{z}-q^{2}/16$ for an odd integer $q$) such that the product $E=f_{1}f_{2}$ of two linearly independent solutions of (1.1) satisfy $\lambda(E)<\infty$ or even $\lambda(E)=0$ [11, Theorem 5.22]. Generally, we have the following
@@
Let $E(z)$ be the product of two linearly independent solutions $f_{1}$ and $f_{2}$ of the differential equation (1.1). Then
@@
A problem left by their work–known as the Bank–Laine conjecture–is whether $\lambda(E)=\infty$ whenever $\rho(A)\not\in\mathbb{N}$. This conjecture has attracted much interest; see the surveys [7, 12] and references therein. Bergweiler and Eremenko [4, 5] disproved this conjecture by using the method of quasiconformal surgery to construct, for each $\rho\in(1/2,\infty)$, entire functions $A(z)$ such that $\rho(A)=\rho$ and the differential equation (1.1) has two linearly independent solutions $f_{1},f_{2}$ such that $\lambda(E)=\rho(E)<\infty$ and one of $f_{1}$ and $f_{2}$ has no zeros. We refer to [1, 13] for the basic results of quasiconformal mapping theory.
An entire function $E(z)$ is called a Bank–Laine function if $E(z)=0$ implies that $E^{\prime}(z)\in\{-1,1\}$. In particular, $E(z)$ is called special if $E(z)=0$ implies that $E^{\prime}(z)=1$. For two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1), if the Wronskian determinant $W(f_{1},f_{2})=f_{1}f_{2}^{\prime}-f_{1}^{\prime}f_{2}=1$, then we say that $f_{1}$ and $f_{2}$ are normalized. A Bank–Laine function $E$ can be written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation (1.1); see [11, Proposition 6.4]. Most study on the Bank–Laine conjecture has concentrated on constructing Bank–Laine functions with certain properties. All Bank–Laine functions have order at least $1$ by Theorem 1.1 and the inequality (1.2). Bergweiler and Eremenko’s construction in [4] shows for the first time that there exists a special Bank–Laine function $E$ such that $\rho(E)$ is not an integer order or half-integer order.
@@
where the expression $S(F)$ is called the Schwarzian derivative of $F$. The Schwarzian derivative is invariant under the linear fractional transformation of $F$. Then $F$ is a locally univalent meromorphic function and a normalized pair $(f_{1},f_{2})$ of solutions of (1.1) can be recovered from $F$ by the formulas $f_{1}^{2}=1/F^{\prime}$ and $f_{2}^{2}=F^{2}/F^{\prime}$. The quotient $F/F^{\prime}=f_{1}f_{2}=E$ is then a Bank–Laine function and all Bank–Laine functions arise in this way. Zeros of a Bank–Laine function $E$ are zeros and poles of $F$ respectively, and $E$ is special if and only if $F$ is entire. By gluing different locally univalent meromorphic functions together using quasiconformal surgery, Bergweiler and Eremenko then recover the desired coefficients $A(z)$ from equation (1.3). In [18], the present author completed Bergweiler and Eremenko’s construction in [5] and, in particular, prove the following
@@
Since $A_{j}$ are all positive numbers, $g^{\prime}_{m,n}(z)\not=0$ for all $z\in\mathbb{C}$ and $g_{m,n}$ is increasing on $\mathbb{R}$, and satisfies $g_{m,n}(x)\to 1$ as $x\to-\infty$ as well as $g_{m,n}(x)\to\infty$ as $x\to\infty$. We also note from the recursive formulas for $A_{i}$ that if $n\geq 1$, then $A_{1}+A_{2}+\cdots+A_{m}<1$ for all $m\geq 1$, which implies that $g_{m,n}$ has no poles on the imaginary axis. If $n=0$ and $m\geq 1$, it is possible that $g_{m,n}$ has some poles on the imaginary axis. Bergweiler and Eremenko’s glued the functions $g_{m,n}$ in (1.5) with $m=0$ and two or infinitely many different even numbers $2n$. Our construction in [18] considered the generic choice of $m$ and $n$. For the underlying ideas behind the gluing, see [5, Section 2].
To prove Theorem 1.2, for each number $\rho\in(1/2,1)$, we let $\gamma=1/(2\rho-1)$. In [18], the author actually proved that, for every $\delta\in[0,1]$, there exists an entire function $A(z)$ of order $\rho(A)=\rho$ such that the differential equation (1.1) has two linearly independent solutions $f_{1}$ and $f_{2}$ such that $E=f_{1}f_{2}$ satisfies $\lambda(E)=\rho\gamma$ and $f_{1}$ and $f_{2}$ satisfy $\lambda(f_{1})=\delta\rho\gamma$ and $\lambda(f_{2})=\rho\gamma$, respectively. In particular, when $\delta=0$ the function $f_{1}$ can be chosen to have no zeros so that the corresponding Bank–Laine function $E$ is special. Theorem 1.2 completes the construction in [5, Corollary 1.1]. As pointed out in [18], we can extend the results to the case that $\rho\in(n/2,n)$ for any positive integer $n$ by following exactly the same process as in [5, Section 5], thus completing the construction in [5, Corollary 1.2]. Unfortunately, the proof of Theorem 1.2 in [18] is not complete because the proof for the two lemmas in [18, Lemma 2.1 and Lemma 2.2] contains gaps in both of the two cases $\gamma=1$ and $\gamma>1$. One of the purposes of this paper is to fill these gaps. The case $\gamma=1$ will be included in Theorem 1.4 below and the corrected versions of the two lemmas will be provided in Section 4.
@@
for two positive constants $c_{1}$ and $c_{2}$ and all large $r$. It seems that the method in [18] does not allow to remove the term $(\log r)^{-2}$ in (1.7). For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, we say that $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $S_{1}(r)$ and $S_{2}(r)$ satisfy $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$. As a complement to Theorem 1.2, we shall first consider gluing two locally univalent meromorphic functions $g_{m,n}$ in (1.5) along the real axis as in [4] and prove the following
@@
Let $M$ and $N$ be two positive integers. Then there exists an entire function $A(z)$ of order
@@
such that the Bank–Laine function $E=f_{1}f_{2}$ for two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(E)=\rho(E)=\rho$ and the number of zeros of $f_{1}$ and $f_{2}$ satisfy $n(r,0,f_{1})\asymp n(r,0,f_{2})$.
By the proof of Theorem 1.3, we can of course choose $M$ and $N$ suitably so that $f_{1}$ has no zeros and thus the corresponding Bank–Laine function $E$ is special. We may even choose $M$ and $N$ suitably so that $f_{2}$ has no zeros.
By looking at the results in Theorem 1.1, Theorem 1.2 and Theorem 1.3, the existence of Bank–Laine functions in the case $\rho(A)$ is an integer is still not clear. Langley [10] pointed out, using Lagrange interpolation, we can make Bank–Laine functions $E=Pe^{Q}$ with $P,Q$ polynomials. However, it is not known if there exist Bank–Laine functions $E(z)$ such that $\lambda(E)=\lambda<\rho(E)=n$ for any preassigned number $\lambda$ and positive integer $n$. It seems that the method of quasiconformal surgery in [5, Subsection 5.3] is invalid for this case, as will be explained in Section 2. In this paper we shall describe a new method of quasiconformal surgery to prove the following
@@
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. Then there exists an entire function $A(z)$ of order $n$ such that the Bank–Laine function $E=f_{1}f_{2}$ for two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$.
@@
As in [18, Theorem 3], we can describe the asymptotic behaviors of the Bank–Laine function $E(z)$ in Theorem 1.3 and also the corresponding coefficient function $A(z)$ of the second order differential equation (1.1) in the case $n=1$. We have the following
@@
where $l$ is an integer such that $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$, while,
@@
The proof for Theorem 1.5 is only a slight modification of that in [18, Theorem 3] for the case $\rho=\sigma=1$ and thus will be omitted. The proof is based on the fact that the Schwarzian derivative $S(F)$ can be factored as $S(F)=B(F/F^{\prime})/2$, where
@@
is a product of two normalized solutions of (1.1). The results in [18, Theorem 3] shows that the functions $A$ and $E$ constructed in Theorem 1.3 are actually of completely regular growth (in the sense of Levin and Pfluger). This is obvious for $E(z)$ and $A(z)$ in the case $\lambda_{1}\leq\lambda_{2}<n$. By similar calculations as in the proof of [18, Theorem 3] with $\rho=\sigma=1$ together with the comments on the extension to the general case in Subsection 3.4, the classical Hadamard factorization theorem actually implies that $E(z)=[H_{1}(z)/H_{2}(z)]e^{p(z)}$, where $H_{1}(z)$ and $H_{2}(z)$ are the canonical products formed by the zeros and poles of $E(z)$ respectively and $p(z)$ is a polynomial of degree $n$ of the form $p(z)=z^{n}+\cdots$.
@@
Let $\rho\in(1/2,\infty)$ and $\lambda\in[0,\infty)$. Then there exists an entire function $A(z)$ of order $\rho$ such that the Bank–Laine function $E=f_{1}f_{2}$ for two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda$ and $\log n(r,0,f_{2})\asymp r^{\rho}$.Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
To prove Theorem 1.6, we not only glue the functions $g_{m,n}$ as in the proof of Theorem 1.2 but also glue the two functions $\hat{g}_{m,n}=(g_{m,n}+1)/2$ and $\hat{g}_{m,n}=g_{m,n}$ along horizontal lines for certain pair $(m,2n)$ of two nonnegative integers $m$ and $n$. If we fix $\hat{g}_{m_{1},n_{1}}=g_{m_{1},n_{1}}$ for one pair $(m_{1},2n_{1})$ of two nonnegative integers $m_{1}$ and $n_{1}$ in the lower half-plane and $\check{g}_{m_{2},n_{2}}=(g_{m_{2},n_{2}}+1)/2$ for another pair $(m_{2},2n_{2})$ of two nonnegative integers $m_{2}$ and $n_{2}$ in the upper half-plane, then we may follow the same process as in the proof of Theorem 1.3 to prove that: for a dense set of orders $\rho\in(1,\infty)$, there exists an entire function $A(z)$ of order $\rho(A)=\rho$ such that the differential equation (1.1) has two linearly independent solutions $f_{1}$ and $f_{2}$ such that $\lambda(f_{1})=\rho$ and $\lambda(f_{2})=\infty$ and no other solution $f_{3}$ linearly independent from $f_{1}$ such that $\lambda(f_{3})<\infty$.
By the proof of Theorem 1.6, the constructed entire function $A(z)$ and the corresponding Bank–Laine function $E$ inherit all the properties of those in Theorem 1.2 in the lower half-plane in the case $\rho\in(1/2,1)$, including the asymptotic behaviors as in [18, Theorem 3]. In fact, by checking the proof for [18, Theorem 3] together with the definition $\hat{g}_{m,n}=(g_{m,n}+1)/2$, we easily find that these asymptotic behaviors also hold for $z=re^{i\theta}$ such that $|\theta|<(1-\varepsilon)\pi/(2\sigma)$ with $\sigma=\rho/(2\rho-1)$. However, in the region $\{z=re^{i\theta}:0\leq\theta\leq\pi/(2\rho)\}$ of the plane, the function $E$ has much complicated behavior. There must be an infinite sequence $(z_{n})$, $z_{n}=r_{n}e^{i\theta_{n}}$, $0\leq\theta_{n}\leq\pi/(2\rho)$ and $r_{n}\to\infty$, such that $\log|E(r_{n}e^{i\theta_{n}})|\geq cr_{n}^{\rho}$ for some positive constant $c$ and all large $r_{n}$.

### v3

@@ Abstract.
A Bank–Laine function $E$ is written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation $f^{\prime\prime}+Af=0$, where $A$ is an entire function. In this paper, we first complete the construction of Bank–Laine functions by Bergweiler and Eremenko and then show the existence of entire functions $A$ for which the associated Bank–Laine functions $E=f_{1}f_{2}$ have preassigned number of zeros of three types:
@@
    for every positive integer $n\in\mathbb{N}$ and every two numbers $\lambda_{1},\lambda_{2}\in[0,n]$ such that $\lambda_{1}\leq\lambda_{2}$, there exists an entire function $A$ of order $n$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$;
@@
    for every number $\rho\in(1/2,\infty)$ and every number $\lambda\in[0,\infty)$, there exists an entire function $A$ of order $\rho$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda$ and $\lambda(f_{2})=\infty$ and $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other $f_{3}$ linearly independent from $f_{1}$;
@@
    for every positive integer $n\in\mathbb{N}$ and every two numbers $\lambda_{1}\in[0,n]$ and $\lambda_{2}\in[0,\infty]$, there exists an entire function $A$ of order $n$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\max\{\lambda_{1},\lambda_{2}\}\leq\rho(E)=\infty$ and $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other $f_{3}$ linearly independent from $f_{1}$.
The constructions for the three types of Bank–Laine functions require new developments of the method of quasiconformal surgery by Bergweiler and Eremenko.
@@ 1. Introduction
(1.1) $f^{\prime\prime}+Af=0,$
where $A$ is an entire function, under the framework of Nevanlinna theory. See [10] for the standard notation and fundamental results of Nevanlinna theory. Recall that, for a nonconstant entire function $f$, the order $\rho(f)$ of $f$ is defined as
@@
where $m(z)$ is the multiplicity of the zero $z$ and $n(r,0,f)=\sum_{|z|\leq r}m(z)$ denotes the number of zeros of $f$ in the disc $\{z:|z|\leq r\}$. It is elementary that all nontrivial solutions of equation (1.1) are entire functions. When $A$ is a transcendental entire function of finite order, an application of the lemma on the logarithmic derivative [10, Section 2.2] easily yields that all nontrivial solutions of (1.1) satisfy $\rho(f)=\infty$. However, there are examples of $A$ (say $A(z)=e^{z}-q^{2}/16$ for an odd integer $q$) such that the product $E=f_{1}f_{2}$ of two linearly independent solutions of (1.1) satisfies $\lambda(E)<\infty$ or even $\lambda(E)=0$ [12, Theorem 5.22]. Generally, we have the following
@@
Let $E$ be the product of two linearly independent solutions $f_{1}$ and $f_{2}$ of the differential equation (1.1). Then
@@
A problem left by their work–known as the Bank–Laine conjecture–is whether $\lambda(E)=\infty$ whenever $\rho(A)\not\in\mathbb{N}$. This conjecture has attracted much interest; see the surveys [8, 13] and references therein. Bergweiler and Eremenko [4, 5] disproved this conjecture by using the method of quasiconformal surgery to construct, for each $\rho\in(1/2,\infty)$, entire function $A$ such that $\rho(A)=\rho$ and the differential equation (1.1) has two linearly independent solutions $f_{1},f_{2}$ such that $\lambda(E)=\rho(E)<\infty$ and one of $f_{1}$ and $f_{2}$ has no zeros. We refer to [1, 14] for the basic results of quasiconformal mapping theory.
An entire function $E$ is called a Bank–Laine function if $E(z)=0$ implies that $E^{\prime}(z)\in\{-1,1\}$. In particular, $E$ is called special if $E(z)=0$ implies that $E^{\prime}(z)=1$. For two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1), if the Wronskian determinant $W(f_{1},f_{2})=f_{1}f_{2}^{\prime}-f_{1}^{\prime}f_{2}=1$, then we say that $f_{1}$ and $f_{2}$ are normalized. A Bank–Laine function $E$ can be written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation (1.1); see [12, Proposition 6.4]. Most study on the Bank–Laine conjecture has concentrated on constructing Bank–Laine functions with certain properties. All Bank–Laine functions have order at least $1$ by Theorem 1.1 and the inequality (1.2). Bergweiler and Eremenko’s construction in [4] shows for the first time that there exists a special Bank–Laine function $E$ such that $\rho(E)$ is not an integer order or half-integer order.
@@
where the expression $S(F)$ is called the Schwarzian derivative of $F$. Then $F$ is a locally univalent meromorphic function and the Schwarzian derivative is invariant under the linear fractional transformation of $F$. This implies a one-to-one correspondence between the equivalence class of locally univalent meromorphic functions and entire function. On the other hand, a normalized pair $(f_{1},f_{2})$ of solutions of (1.1) can be recovered from $F$ by the formulas $f_{1}^{2}=1/F^{\prime}$ and $f_{2}^{2}=F^{2}/F^{\prime}$. The quotient $F/F^{\prime}=f_{1}f_{2}=E$ is then a Bank–Laine function and all Bank–Laine functions arise in this way. Zeros of a Bank–Laine function $E$ are zeros and poles of $F$ respectively, and $E$ is special if and only if $F$ is entire. By gluing different locally univalent meromorphic functions together using quasiconformal surgery, Bergweiler and Eremenko then recover the desired $F$ and also the corresponding coefficient $A$ from equation (1.3). In [19], the present author completed Bergweiler and Eremenko’s construction in [5] and, in particular, prove the following
@@
Since $A_{j}$ are all positive numbers, $g^{\prime}_{m,n}(z)\not=0$ for all $z\in\mathbb{C}$ and $g_{m,n}$ is increasing on $\mathbb{R}$, and satisfies $g_{m,n}(x)\to 1$ as $x\to-\infty$ as well as $g_{m,n}(x)\to\infty$ as $x\to\infty$. By the recursive formulas for $A_{i}$, we see that if $n\geq 1$, then $A_{1}+A_{2}+\cdots+A_{m}<1$ for all $m\geq 1$, which implies that $g_{m,n}$ has no poles on the imaginary axis. If $n=0$ and $m\geq 1$, it is possible that $g_{m,n}$ has some poles on the imaginary axis. Bergweiler and Eremenko’s glued the functions $g_{m,n}$ in (1.5) with $m=0$ and two or infinitely many different even numbers $2n$. Our construction in [19] considered the generic choice of $m$ and $n$. For the underlying ideas behind the gluing, see [5, Section 2].
To prove Theorem 1.2, for each number $\rho\in(1/2,1)$, we let $\gamma=1/(2\rho-1)$. In [19], the author actually proved that, for every $\delta\in[0,1]$, there exists an entire function $A$ of order $\rho(A)=\rho$ such that the differential equation (1.1) has two linearly independent solutions $f_{1}$ and $f_{2}$ such that $E=f_{1}f_{2}$ satisfies $\lambda(E)=\rho\gamma$ and $f_{1}$ and $f_{2}$ satisfy $\lambda(f_{1})=\delta\rho\gamma$ and $\lambda(f_{2})=\rho\gamma$, respectively. In particular, when $\delta=0$ the function $f_{1}$ can be chosen to have no zeros so that the associated Bank–Laine function $E$ is special. Theorem 1.2 completes the construction in [5, Corollary 1.1]. As pointed out in [19], we can extend the results to the case that $\rho\in(n/2,n)$ for any positive integer $n$ by following exactly the same process as in [5, Section 5], thus completing the construction in [5, Corollary 1.2]; see also the comments in Subsection 3.5 and Subsection 4.3. Unfortunately, the proof of Theorem 1.2 in [19] is not complete because the proof for the two lemmas in [19, Lemma 2.1 and Lemma 2.2] contains gaps in both of the two cases $\gamma=1$ and $\gamma>1$. One of the purposes of this paper is to fill these gaps. The case $\gamma=1$ will be included in Theorem 1.4 below and the corrected versions of the two lemmas will be provided in Section 4.
@@
for two positive constants $c_{1}$ and $c_{2}$ and all large $r$. Thus $f_{1}$ has less number of zeros than $f_{2}$. It seems that the method in [19] does not allow to remove the term $(\log r)^{-2}$ in (1.7). For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, we say that $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $S_{1}(r)$ and $S_{2}(r)$ satisfy $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$. As a complement to Theorem 1.2, we shall first consider gluing two locally univalent meromorphic functions $g_{m,n}$ in (1.5) along the real axis as in [4] and prove the following
@@
Let $M$ and $N$ be two positive integers such that $M\not=N$. Then there exists an entire function $A$ of order
@@
such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(E)=\rho(E)=\rho$ and the number of zeros of $f_{1}$ and $f_{2}$ satisfy $n(r,0,f_{1})\asymp n(r,0,f_{2})$.
By the proof of Theorem 1.3, we know that the zeros of $f_{1}$ and $f_{2}$ actually satisfy $n(r,0,f_{i})\asymp r^{1/\text{Re}(\mu)}(\log r)$, $i=1,2$. Moreover, if we assume that $M=N$, then the term $\log r$ here can be removed. We remark that we can choose $M$ and $N$ suitably so that $f_{1}$ has no zeros and thus the corresponding Bank–Laine function $E$ is special. We may even choose $M$ and $N$ suitably so that $f_{2}$ has no zeros.
By looking at the results in Theorem 1.1, Theorem 1.2 and Theorem 1.3, the existence of Bank–Laine functions in the case $\rho(A)$ is an integer is still not clear. Langley [11] pointed out, using Lagrange interpolation, we can make Bank–Laine functions $E=Pe^{Q}$ with $P,Q$ being polynomials. However, it is not known if there exist Bank–Laine functions $E$ such that $\lambda(E)=\lambda<\rho(E)=n$ for any preassigned number $\lambda$ and positive integer $n$. Moreover, all the constructions now yield Bank–Laine functions such that $\rho(E)=\rho(A)$, but it is not known if there exist Bank–Laine functions such that $\lambda(E)>\rho(A)$. It seems that the method of quasiconformal surgery in [5, Subsection 5.3] is invalid when dealing with these two cases, as will be explained in Section 3. In this paper we shall describe a new method of quasiconformal surgery for constructions of such kind of Bank–Laine functions. Concerning the first question, we shall prove the following
@@
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. Then there exists an entire function $A$ of order $n$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$.
@@
As in [19, Theorem 3], we can describe the asymptotic behaviors of the Bank–Laine function $E$ in Theorem 1.3 and also the corresponding coefficient function $A$ of the second order differential equation (1.1) in the case $n=1$. We have the following
@@
where $l$ is an integer such that $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$, while
@@
The proof for Theorem 1.5 is only a slight modification of that in [19, Theorem 3] for the case $\rho=\sigma=1$ and thus will be omitted. The proof is based on the fact that the Schwarzian derivative $S(F)$ can be factored as $S(F)=B(F/F^{\prime})/2$, where
@@
is a product of two normalized solutions of (1.1). The results in [19, Theorem 3] shows that the functions $A$ and $E$ constructed in Theorem 1.3 are actually of completely regular growth (in the sense of Levin and Pfluger). This is obvious for $E$ and $A$ when $\lambda_{1}\leq\lambda_{2}<n$. In fact, by the classical Hadamard factorization theorem, Theorem 1.5 together with its extension to the general case in Subsection 3.4 implies that $E(z)=H(z)e^{p(z)}$, where $H$ is the canonical product formed by the zeros of $E$ and $p$ is a polynomial of degree $n$ of the form $p(z)=z^{n}+\cdots$.
@@
Let $\rho\in(1/2,\infty)$ and $\lambda\in[0,\infty)$. Then there exists an entire function $A$ of order $\rho$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda$ and $\log n(r,0,f_{2})\asymp r^{\rho}$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
To prove Theorem 1.6, we not only glue the functions $g_{m,n}$ as in the proof of Theorem 1.2 but also glue the two functions $\hat{g}_{m,n}=(g_{m,n}+1)/2$ and $\hat{g}_{m,n}=g_{m,n}$ along horizontal lines for certain pair $(m,2n)$ of two nonnegative integers $m$ and $n$. If we fix $\check{g}_{m_{1},n_{1}}=g_{m_{1},n_{1}}$ for one pair $(m_{1},2n_{1})$ of two nonnegative integers $m_{1}$ and $n_{1}$ in the lower half-plane and $\hat{g}_{m_{2},n_{2}}=(g_{m_{2},n_{2}}+1)/2$ for another pair $(m_{2},2n_{2})$ of two nonnegative integers $m_{2}$ and $n_{2}$ in the upper half-plane, then we may follow the same process as in the proof of Theorem 1.3 to prove: for a dense set of orders $\rho\in(1,\infty)$, there exists an entire function $A$ of order $\rho(A)=\rho$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})<\infty$ and $\lambda(f_{2})=\infty$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
By the proof of Theorem 1.6, the constructed entire function $A$ and the corresponding Bank–Laine function $E$ inherit all the properties of those in Theorem 1.2 in the lower half-plane in the case $\rho\in(1/2,1)$, including the asymptotic behaviors as in [19, Theorem 3]. In fact, by checking the proof for [19, Theorem 3] together with the definition $\hat{g}_{m,n}=(g_{m,n}+1)/2$, we easily find that these asymptotic behaviors also hold for $z=re^{i\theta}$ such that $|\theta|<(1-\varepsilon)\pi/(2\sigma)$ with $\sigma=\rho/(2\rho-1)$. However, in the region $\{z=re^{i\theta}:0\leq\theta\leq\pi/(2\rho)\}$ of the plane, the function $E$ has much complicated behavior. There must be an infinite sequence $(z_{n})$, $z_{n}=r_{n}e^{i\theta_{n}}$, $0\leq\theta_{n}\leq\pi/(2\rho)$ and $r_{n}\to\infty$, such that $\log|E(r_{n}e^{i\theta_{n}})|\geq cr_{n}^{\rho}$ for some positive constant $c$ and all large $r_{n}$.